# A variational formulation of the free energy of mixed quantum-classical systems: coupling classical and electronic density functional theories

### v4

@@ II.3.2 Rewriting of the functional $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$
Because $\boldsymbol{p}$ does not depend on the quantum variables and that we chose the normalization condition of Eq. (29), we can naturally decompose the entropy as
@@
with $S_{{\text{mm}}}[\boldsymbol{p}]$ being the purely classical dimension-less entropy
@@
$S_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ $=-\text{Tr}\big{\{}\hat{{\boldsymbol{\rho}}}\boldsymbol{p}\log(\hat{{\boldsymbol{\rho}}})\big{\}}$ (39)
@@
It is noteworthy here to highlight that the form of Eq. (45) is quite general as it consists in a priori in potential with a non local kernel ${\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}},{\bf{q}},{\bf{q}}^{\prime})$. This form of interaction includes for instance Lennard-Jones potentials, frozen density Coulomb interaction or density overlap models for Pauli repulsionWheatley and Price (1990).
@@
[…]
$+E_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]-\text{k}_{\text{B}}T\big{(}S_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]+S_{{\text{mm}}}[\boldsymbol{p}]\big{)}\big{\}}.$
Eq. (46) is therefore a variational formulation of $F_{0}$ for a mixed QM/MM system in terms of a minimization over the couple $(\hat{{\boldsymbol{\rho}}},\boldsymbol{p})$, which, to the best of our knowledge, was never formalized before. One can see from Eqs. (38), (41) and (42) that the functionals corresponding to the MM part of the energy and entropy are completely independent of the quantum part of the system, as they depend only on the classical probability $\boldsymbol{p}$. Nevertheless, because of the term $E_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$, the quantum and classical parts are coupled in the minimization process of Eq. (46). Also, one can see that the functionals corresponding to the QM parts of the energy and entropy explicitly depend on the classical MM probability distribution $\boldsymbol{p}$.
Although Eq. (46) provides a variational formulation of the Helmholtz free energy for a QM/MM system, the variables $(\hat{{\boldsymbol{\rho}}},\boldsymbol{p})$ involved in the minimization are nevertheless very high-dimensional. Therefore, the goal of the following sections is to reformulate the variational principle of Eq. (46) using functionals of significantly lower-dimensional densities, i.e. the one-body QM and MM densities.
@@ III One-body density formulation through Levy-Lieb constrained search formulation
An important consequence of the one-body nature of both $V^{\text{ext}}_{{\text{mm}}}({Q})$ and $v^{\text{ext}}_{{\text{qm}}}({q})$, along with the permutational symmetry of both $\boldsymbol{p}$ and $\hat{{\boldsymbol{\rho}}}$ is that the corresponding energy terms $E_{\text{mm}}^{\text{ext}}[\boldsymbol{p}]$ and $E_{\text{qm}}^{\text{ext}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ in Eqs. (42) and (44) can be computed using one-body objects. Therefore, by applying the Levy-Lieb constrained search, we can reformulate the variational problem of Eq. (46) in terms of a functional that depends only on the one-body quantum and classical densities, rather than the full $N_{{\text{mm}}}+N_{{\text{qm}}}$ variables.
@@ III.1 Classical and quantum one-body density and the Levy-Lieb constrained search
We naturally introduce the one-body classical density $n({Q})$ as the trace over all quantum variables and all classical variables but one of the product $\hat{{\boldsymbol{\rho}}}({\bf{q}},{\bf{Q}})\boldsymbol{p}({\bf{Q}})$, which, thanks to the normalization of Eq. (29), is simply
@@
It is noteworthy from Eq. (48) that, because of the ${\bf{Q}}$ dependence in $\hat{{\boldsymbol{\rho}}}({\bf{Q}})$, the classical $N$-body probability $\boldsymbol{p}({\bf{Q}})$ contributes to the one-body quantum density $\rho({q})$, whereas from Eq. (47) it is clear that the quantum $N$-body density matrix $\hat{{\boldsymbol{\rho}}}({\bf{Q}})$ does not contribute to the classical one-body density.
By using permutation symmetry of both $\hat{{\boldsymbol{\rho}}}({\bf{Q}})$ and $\boldsymbol{p}({\bf{Q}})$ with respect to exchange of either QM or MM variables, we can then write the two terms arising from the external potential as functionals of the QM and MM one-body densities
@@
As the contributions coming from the external potentials can be computed directly with a couple of one-body object $(\rho,n)$, we will now introduce a universal functional of $(\rho,n)$ which contains all necessary information to compute the Gibbs free energy. This functional is built using the Levy-Lieb constrained search formalism as in eDFT. First, we introduce the intrinsic Gibbs free energy functional of the $N$-body objects
$f[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ $=E_{{\text{mm}}}[\boldsymbol{p}]+E_{\text{qm}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]+E_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]-\text{k}_{\text{B}}T\mathcal{S}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}].$ (51)
such that we can then rewrite the Gibbs free energy as
@@
where the notation $(\hat{{\boldsymbol{\rho}}},\boldsymbol{p})\rightarrow(\rho,n)$ means that the search is done over the set of couple of $N$-body objects $(\hat{{\boldsymbol{\rho}}},\boldsymbol{p})$ having the couple of 1-body object given by $(\rho,n)$. Therefore, by remembering that the external parts depends only on the one-body densities (see Eqs. (49) and (50)), one can rewrite Eq. (53) as
@@
Therefore, we can introduce the universal Levy-Lieb functional of the one-body classical and quantum density as
@@
$F_{0}=\min_{(\rho,n)}\big{(}(V^{\text{ext}}_{{\text{mm}}}|n)+(v^{\text{ext}}_{{\text{qm}}}|\rho)+\mathcal{F}[\rho,n]\big{)},$ (56)
which achieves to prove that we can obtain $F_{0}$ as a minimization over only one-body densities.
@@ III.2 Explicit coupling in terms of eDFT and cDFT
Although Eq. (56) allows us to obtain the Gibbs free energy only in terms of the one-body densities $(\rho,n)$, such a functional remains unknown and the explicit connection with well known universal quantum and classical density functionals is yet unclear. This paragraph is concerned with a rewriting of $\mathcal{F}[\rho,n]$ such that the functional of MerminMermin (1965) and EvansEvans (1979) explicitly appear, together with a new QM/MM correlation functional.
@@
$\mathcal{F}_{\text{qm}}[\rho]=\min_{\hat{{\boldsymbol{\rho}}}\rightarrow\rho}\text{Tr}\{\hat{{\boldsymbol{\rho}}}\big{(}\hat{\bf t}+\hat{\bf w}_{{\text{qm}}}+\text{k}_{\text{B}}T\log(\hat{{\boldsymbol{\rho}}})\big{)}\},$ (57)
@@
$\mathcal{F}_{\text{mm}}[n]=\min_{\boldsymbol{p}\rightarrow n}\text{Tr}\{\boldsymbol{p}\big{(}{\bf W}_{{\text{mm}}}+\text{k}_{\text{B}}T\log(\boldsymbol{p})\big{)}\},$ (58)
@@
which will be the QM/MM equivalent of the Hartree-term in eDFT, and where $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ is the QM/MM correlation functional defined of course as the complementary functional
@@
[…]
@@
where we can explicitly recognize the purely quantum and purely classical free-energy functionals, together with the QM/MM coupling expressed as the sum of the mean-field and correlation terms. If $\delta\mathcal{F}_{{\text{qm}}}$ is ignored, we recover the mean-field coupling between the eDFT and cDFT functionals that is generally used in the literaturePetrosyan, Rigos, and Arias (2005); Petrosyan et al. (2007); Tang et al. (2020); Jeanmairet, Levesque, and Borgis (2020); Labat, Giner, and Jeanmairet (2024).
@@ III.3 Applications to solvation problems
We now assume that, in the absence of the solvent, the typical electronic gap is large compared to the thermal agitation $\text{k}_{\text{B}}T$, allowing us to neglect the contribution of electronic entropy in $\mathcal{F}_{\text{qm}}[\rho]$ (see Eq. (57)). This is equivalent to considering only the ground state of the electronic quantum system. To illustrate the validity of this approximation for systems at room temperature ($T=300\,K$), we note that the Boltzmann weight ratio between the ground state and an excited state with a gap of 0.25 eV ($\approx 0.01\,\text{Ha}$) is approximately $10^{-5}$. For a gap of 1.63 eV ($\approx 0.06\,\text{Ha}$), which marks the onset of the visible spectrum, this ratio drops dramatically to $10^{-28}$. Since most chemical systems near their equilibrium geometry exhibit electronic gaps of several eV, this approximation is well-suited for typical room-temperature conditions. As a consequence, we can take the limit $T\rightarrow 0$ in the definition of the Mermin universal functional (see Eq. (57)) to recover the usual universal Levy-Lieb electronic density functional, namely
$\mathcal{F}_{\text{qm}}[\rho]\approx\min_{\Psi\rightarrow\rho}{\langle\Psi|}\hat{\bf t}+\hat{\bf w}_{{\text{qm}}}{|\Psi\rangle}.$ (63)
Regarding solvent effects, we focus on cases where the solvent-solute interaction is not too strong, meaning there is no covalent bonding between the solute and solvent species. Under these conditions, the previous $T=0~{}K$ assumption for $\mathcal{F}_{\text{qm}}[\rho]$ remains reasonable, as the electronic structure of the solute is not significantly altered by the presence of the solvent.
In previous work about mixed electronic-classical density functional theory to study QM/MM systems, the classical fluid was described in the Grand Canonical ensemble. This is because classical DFT is naturally formulated in this ensemble. In practice this means that the number of classical particle is now allowed to fluctuate by exchanging with a reservoir of fixed chemical potential $\mu$. In this ensemble, Eq. (62)
@@
$+\varepsilon_{{\text{qm}}}^{{\text{mm}}}[\rho,n]+\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]\big{\}}$
where $\Omega_{0}$ is the grand potential, that is the state function in the grand canonical ensemble, and $\mathcal{F}_{\text{mm}}$ is now the grand canonical universal functional a la Evans.
In Eq. (64) one recognizes the usual quantum Born-Oppenheimer part for the electron (i.e. $(v_{\text{ne}}|\rho)+\mathcal{F}_{\text{qm}}[\rho]$), together with the purely classical part of the solvent (i.e. $\mathcal{F}_{\text{mm}}[n]+(v_{nucl}^{{\text{mm}}}-\mu|n)$), and $\varepsilon_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ accounts for the mean-field electron-solvent interaction, while $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ accounts for solvation correlation effects in the QMMM system.
In practice, Eq. (64) requires some approximations. First, the choice of the functionals describing the quantum system and the liquid considerably influences the quality of the predictions. However, determining the appropriate electronic functional to describe a given solute and the suitable functional for a classical fluid are topics that themselves deserve a thorough review. Therefore, we let this discussion aside to focus on the QM/MM coupling. In Petrosyan et al’s joint DFT Petrosyan, Rigos, and Arias (2005); Petrosyan et al. (2007), the interaction of the solvent with both the solute nuclei and electrons is dealt as a whole and the coupling is described by the use of pseudo potential. Thus Eq. (64) solely contains the quantum functional $\mathcal{F}_{\text{qm}}$, the solvent one $\mathcal{F}_{\text{mm}}$ and $\varepsilon_{{\text{qm}}}^{{\text{mm}}}$. The coupling term is modeled as a dielectric screening due to a non local dielectric function depending on the electron density and a mean-field term between both densities interacting through a Gaussian repulsive potential.
@@
We also proposed to couple wave function theory-based methods, such as Hartree-Fock and selected configuration interaction with cDFTLabat, Giner, and Jeanmairet (2024) to approximate the universal Levy-Lieb functional $\mathcal{F}_{\text{qm}}[\rho]$ of Eq.62.

### v5

@@ II.3.2 Rewriting of the functional $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$
Because the function $\boldsymbol{p}$ does not explicitly depend on the quantum variables and because we chose the normalization conditions of Eq. (32), we can naturally decompose the entropy as
@@
where $S_{{\text{mm}}}[\boldsymbol{p}]$ is the purely classical dimension-less entropy
@@
[…]
@@
It is worth noting that the form of Eq. (48) is quite general, since ${\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}},{\bf{q}},{\bf{q}}^{\prime})$ is a priori non local. This form of interaction includes, for example, Lennard-Jones potentials, frozen density Coulomb interaction or density overlap models for Pauli repulsionWheatley and Price (1990).
@@
$F_{0}=\min_{\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}}\big\{$ $E_{\text{qm}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]+E_{\text{qm}}^{\text{ext}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]+E_{{\text{mm}}}[\boldsymbol{p}]+E_{\text{mm}}^{\text{ext}}[\boldsymbol{p}]$ (49)
$+E_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]-T\big(S_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]+S_{{\text{mm}}}[\boldsymbol{p}]\big)\big\}.$
Eq. (49) therefore provides a variational formulation of $F_{0}$ for a mixed QM/MM system, expressed as a minimization over the couple $(\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p})$. One can see from Eqs. (41), (44) and (45) that the functionals corresponding to the MM part of the energy and entropy (i.e. $E_{\text{mm}}^{\text{ext}}[\boldsymbol{p}]$, $E_{{\text{mm}}}[\boldsymbol{p}]$ and $S_{{\text{mm}}}[\boldsymbol{p}]$) are, per se, completely independent of the quantum part of the system. They only depend on the classical probability $\boldsymbol{p}$. Nevertheless, the quantum and classical parts remain coupled in the minimization of Eq. (49) through the coupling functional $E_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$. In contrast, the QM parts energy and entropy functionals explicitly depend on the classical MM probability distribution $\boldsymbol{p}$.
Although Eq. (49) provides a variational formulation of the Helmholtz free energy for a QM/MM system, the minimization variables $(\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p})$ remain very high-dimensional because they explicitly describe every individual particle in the system. The objective of the following section is to reformulate the variational principle of Eq. (49) using functionals depending on much lower-dimensional objects, i.e. the one-body QM and MM densities.
@@ III One-body density formulation through Levy-Lieb constrained search formulation
Starting from the variational formulation of the Helmholtz free energy for the QM/MM system in terms of the $N_{{\text{qm}}}+N_{{\text{mm}}}$ quantities in Eq. (49), we want to obtain a one-body DFT formulation. As in classical and quantum DFT, the key idea is to split the total energy into an external part and an intrinsic part. More precisely, because of the one-body nature of both classical and quantum external potentials $V^{\text{ext}}_{{\text{mm}}}({Q})$ and $v^{\text{ext}}_{{\text{qm}}}({q})$, and the permutational symmetry of both $\boldsymbol{p}$ and $\hat{{\boldsymbol{\rho}}\,}$, the corresponding energy terms $E_{\text{mm}}^{\text{ext}}[\boldsymbol{p}]$ and $E_{\text{qm}}^{\text{ext}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ in Eqs. (45) and (47) can be expressed in terms of the one-body densities. Then, by applying the Levy-Lieb constrained search, we can also reformulate the intrinsic part of the energy as a functional of the one-body quantum and classical densities. Therefore, the variational problem of Eq. (49) is directly expressed as a functional of the one-body QM and MM densities rather than in terms of full $N_{{\text{mm}}}+N_{{\text{qm}}}$ variables quantities.
@@ III.1 Classical and quantum one-body density and the Levy-Lieb constrained search
The classical one-body density is naturally defined as the trace over all quantum variables and all classical variables but one of the product $\hat{{\boldsymbol{\rho}}\,}({\bf{q}},{\bf{Q}})\boldsymbol{p}({\bf{Q}})$, which, thanks to the normalization of Eq. (32), is simply
@@
Thanks to the permutational symmetry of both $\hat{{\boldsymbol{\rho}}\,}({\bf{Q}})$ and $\boldsymbol{p}({\bf{Q}})$ with respect to the exchange of either QM or MM coordinates, the two external functionals can be rewritten as functionals of the QM and MM one-body densities, respectively
@@
Since the contributions arising from the external potentials can be directly expressed in terms of the one-body densities $(\rho,n)$, we now introduce a universal functional of $(\rho,n)$ which contains all the information required to compute the free energy. This functional is built using the Levy-Lieb constrained search formalism, as in eDFT. First, we introduce the intrinsic Helmholtz free energy functional of the $N$-body densities
$f[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ $=E_{{\text{mm}}}[\boldsymbol{p}]+E_{\text{qm}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]+E_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]-T\mathcal{S}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}].$ (54)
such that we can then rewrite the Helmholtz free energy as
@@
where the notation $(\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p})\rightarrow(\rho,n)$ means that the search is performed over the set of couple of $N_{{\text{qm}}}+N_{{\text{mm}}}$-body densities $(\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p})$ that yield the 1-body densities $(\rho,n)$. Therefore, by remembering that the external parts depends only on the one-body densities (see Eqs. (52) and (53)), one can rewrite Eq. (56) as
@@
We have just introduced the universal Levy-Lieb functional of the one-body classical and quantum density,
@@
$F_{0}=\min_{(\rho,n)}\big((V^{\text{ext}}_{{\text{mm}}}|n)+(v^{\text{ext}}_{{\text{qm}}}|\rho)+\mathcal{F}[\rho,n]\big).$ (59)
This provides a constructive proof of the existence of a variational principle for the Helmholtz Free energy, $F_{0}$, as a functional of one-body densities.
@@ III.2 Explicit coupling in terms of eDFT and cDFT
Although Eq. (59) expresses the Helmholtz free energy only in terms of the one-body densities $(\rho,n)$, the functional itself remains unknown in practice. Thus, the explicit connection with well known universal quantum and classical density functionals is still unclear. This paragraph focuses on rewriting of $\mathcal{F}[\rho,n]$ so that the functionals of MerminMermin (1965) and EvansEvans (1979) explicitly appear, along with a new QM/MM correlation functional.
@@
[…]
@@
$\mathcal{F}_{\text{mm}}[n]=\min_{\boldsymbol{p}\rightarrow n}\text{Tr}\{\boldsymbol{p}\big({\bf W}_{{\text{mm}}}+T\log(\boldsymbol{p})\big)\},$ (61)
@@
which is the QM/MM analogue of the Hartree-term in eDFT, and where $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ is the QM/MM correlation functional, i.e, the complementary functional
@@
$F_{0}=\min_{(\rho,n)}\big\{$ $(v^{\text{ext}}_{{\text{qm}}}|\rho)+\mathcal{F}_{\text{qm}}[\rho]+\mathcal{F}_{\text{mm}}[n]+(V^{\text{ext}}_{{\text{mm}}}|n)$ (65)
@@
where we can explicitly recognize the purely quantum and purely classical free-energy functionals, together with the QM/MM coupling expressed as the sum of the mean-field and correlation terms. If $\delta\mathcal{F}_{{\text{qm}}}$ is ignored, we recover the mean-field coupling between the eDFT and cDFT functionals that is generally used in the literaturePetrosyan, Rigos, and Arias (2005); Petrosyan et al. (2007); Tang et al. (2020); Jeanmairet, Levesque, and Borgis (2020); Labat, Giner, and Jeanmairet (2024). We highlight here that, although the outcome of the present derivation is rather natural as it consists of a multi component DFT formalism, the explicit definition of $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ in Eq. (64) can help in future developments of approximations for $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$. For instance, the variational formulation in terms of full-body densities allows the use the adiabatic connection to reformulate the problem, and therefore potentially to import several successful strategies used in eDFT to approximate correlation functionals.
@@ III.3 Applications to solvation problems
We now assume that, in the absence of the solvent, the typical electronic gap is large compared to the thermal agitation $T$, allowing us to neglect the contribution of electronic entropy in $\mathcal{F}_{\text{qm}}[\rho]$ (see Eq. (60)). This is equivalent to considering only the ground state of the electronic quantum system. To illustrate the validity of this approximation for systems at room temperature (T=300KT=300\,K), we note that the Boltzmann weight ratio between the ground state and an excited state with a gap of 0.25 eV ($\approx 0.01\,\text{Ha}$) is approximately $10^{-5}$. For a gap of 1.63 eV ($\approx 0.06\,\text{Ha}$), which marks the onset of the visible spectrum, this ratio drops dramatically to $10^{-28}$. Since most chemical systems near their equilibrium geometry exhibit electronic gaps of several eV, this approximation is well-suited for typical room-temperature conditions. As a consequence, we can take the limit $T\rightarrow 0$ in the definition of the Mermin universal functional (see Eq. (60)) to recover the usual universal Levy-Lieb electronic density functional, namely
@@
Regarding solvent effects, we focus on cases where the solvent-solute interaction is not too strong, meaning there is no covalent bonding between the solute and solvent species. Under these conditions, the previous T=0KT=0~K assumption for $\mathcal{F}_{\text{qm}}[\rho]$ remains reasonable, as the electronic structure of the solute is not significantly altered by the presence of the solvent.
Previous attempts to apply mixed electronic-classical density functional theory to QM/MM systems were carried on in a semi-Grand canonical ensemble, where the number of classical particles is allowed to fluctuate. This choice was motivated by the availability of classical functionals, which are naturally formulated in the grand canonical ensemble. In the semi-Grand canonical ensemble, the state function is the grand potential, $\Omega_{0}$, for which the following variational principle, analogous to Eq. (65), stands
$\Omega_{0}=$ $\min_{(\rho,n)}\big\{\mathcal{F}_{\text{qm}}[\rho]+\Omega_{\text{mm}}[n^{\text{GC}}]+\delta\Omega_{{\text{qm}}}^{\text{mm}}[\rho,n^{\text{GC}}]$ (67)
$+\varepsilon_{{\text{qm}}}^{\text{mm}}[\rho,n^{\text{GC}}]+\left(v^{\text{ext}}_{{\text{qm}}}\mid\rho\right)+\left(V^{\text{ext}}_{{\text{mm}}}\mid n^{\text{GC}}\right)\big\}.$
In Eq. (67), $n^{\text{GC}}$ is the one-body density of classical particles, $\Omega_{\text{mm}}[n^{\text{GC}}]$ is the grand canonical universal functional a la Evans, $V^{\text{ext}}_{{\text{mm}}}$ is the external classical potential, $(v_{\text{ne}}|\rho)+\mathcal{F}_{\text{qm}}[\rho]$) is the usual quantum Born-Oppenheimer electronic functional and $\delta\Omega_{{\text{qm}}}^{\text{mm}}[\rho,n^{\text{GC}}]$ is the unknown QM/MM correlation term. Eq. (67) can be obtained following the same route that was used to derive Eq. (65). Its explicit derivation is in available in S.I. III.
Minimizing Eq. (67) in practice require some approximations. First, the choice of the functionals describing the quantum system and the liquid considerably influences the quality of the predictions. However, determining the appropriate electronic functional for a given solute and the suitable functional for a classical fluid are complex topics that themselves deserve a thorough review. Therefore, we let this discussion aside to focus on the QM/MM coupling and the several strategies adopted in the litterature.
In Petrosyan et al’s joint DFT Petrosyan, Rigos, and Arias (2005); Petrosyan et al. (2007), the interaction of the solvent with both the solute nuclei and electrons is dealt as a whole and the coupling is described by the use of pseudo potential. Thus Eq. (67) solely contains the quantum functional $\mathcal{F}_{\text{qm}}$, the solvent one $\mathcal{F}_{\text{mm}}$ and $\varepsilon_{{\text{qm}}}^{{\text{mm}}}$. The coupling term is modeled as a dielectric screening due to a non local dielectric function depending on the electron density and a mean-field term between both densities interacting through a Gaussian repulsive potential.
@@
We also proposed to couple wave function theory-based methods, such as Hartree-Fock and selected configuration interaction with cDFTLabat, Giner, and Jeanmairet (2024) to approximate the universal Levy-Lieb functional $\mathcal{F}_{\text{qm}}[\rho]$ of Eq. (65).